OpenStudy (anonymous):

A line has a slope of 3 and contains the point (−1, −8). Which equations represent the line? Choose all answers that are correct. A. −3x + y = −5 B. 3x + y = −5 C. (y – 8) = 3(x − 1) D. (y + 8) = 3(x + 1)

OpenStudy (anonymous):

The point slope formula is \[\huge~\rm~y-y_1=m(x-x_1)\] whereas m=slope (x1,y1)

OpenStudy (anonymous):

Can you sub in the vaules?

OpenStudy (anonymous):

A and D?

OpenStudy (anonymous):

@Moe95

OpenStudy (anonymous):

Oh wait thats correct

OpenStudy (anonymous):

You have it correct :)

OpenStudy (anonymous):

Thank You! :)

OpenStudy (anonymous):

:)